# Rate-reliability functions

### v2

@@ IV-B Zero-dimensional output probability sets
Example 12 can be generalized to DMCs with an input power constraint. Given a cost function $\phi:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ (power per symbol), an input power constraint $A$ is an upper bound on the average power that admissible code words have to obey: $\phi^{n}(x^{n})=\frac{1}{n}\sum_{t=1}^{n}\phi(x_{t})\leq A$. From [SPBD:DI_power] we know that the DI capacity of the DMC $W$ under the power constraint $A$ is
$C_{\text{DI}}(W;\phi,A)=\max_{p_{X}\,:\,\mathbb{E}\phi(X)\leq A}H(X),$
after purging the channel of duplicate input symbols: if $W_{x}=W_{x^{\prime}}$, and say $\phi(x^{\prime})\geq\phi(x)$, remove the input with the larger cost ($x^{\prime}$) from $\mathcal{X}$ and keep the other one ($x$); in case of a tie decide arbitrarily. Example 12 above is recovered when the power constraint is irrelevant (if $A\geq\max_{x\in\mathcal{X}}\phi(x)$), in which case the capacity becomes $\max_{p_{X}\in\mathcal{P}(\mathcal{X})}H(X)=\log|\mathcal{X}|=\log|W(\mathcal{X})|$. With the power constraint active, and for positive error exponents, we obtain lower and upper bounds on $R(n)$ as given in Eqs. (25) and (26), with $\log|W(\mathcal{X})|$ replaced by $\max_{p_{X}\,:\,\mathbb{E}\phi(X)\leq A}H(X)$.
@@
IV-C Extension to quantum channels
@@
Here we prove how the main results in Section III can be extended to classical-quantum (cq)-channels and to general quantum channels under the restriction that we only use product states when encoding. We refer the reader to [CDBW:DI_classical, Sect. 6] for all the necessary quantum information tools (alternatively, see [Wilde-book, holevo:stat-structure-book, CK:book2011]) needed for the following, as well as for a brief review on identification via quantum channels [loeber:PhDThesis, AW:StrongConverse, Winter:QCidentification, Winter:QC-ID-randomness, Winter:Review, CDBW:ID-0-sim, CDBW:ID-0-sim:ICC].
@@
For the direct part, we just need to define an informationally complete POVM $T=(T_{y}:y\in\mathcal{Y})$. Then, we can form the concatenated channel $\overline{W}=\mathcal{T}\circ W$ of the cq-channel $W$ followed by the quantum-classical (qc)-channel $\mathcal{T}(\sigma)=\sum_{y}(\operatorname{Tr}\sigma T_{y})\ket{y}\bra{y}$ |, which effectively outputs the probability of each outcome $y$ when the POVM $T$ is used as measurement. Now, since that the channel $\overline{W}$ can be identified with the classical channel $\overline{W}(y|x)=\operatorname{Tr}W_{x}T_{y}$ and $\widehat{\mathcal{X}}:=\overline{W}(\mathcal{X})\subset\mathcal{P}(\mathcal{Y})$ its image of output distributions we can reduce the problem directly to the setup of Theorem 4 to get the following:
@@
For any $10<t<$ 1, and $E(n)>0$, there exists a simultaneous DI code over the cq-channel $W:\mathcal{X}\rightarrow\mathcal{S}(B)$ (and its associated $\overline{W}$ as above) on block length $n$ with error exponents $E_{1}(n)\geq E(n)-\frac{1}{n}$ and $E_{2}(n)\geq E(n)-\frac{3}{n}$, and its rate lower-bounded
$\begin{split}R(n)&\geq(1-t)\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widehat{\mathcal{X}}}\right)\right]\\ &\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right).\qed\end{split}$
@@
With the same arguments, we can also extend Corollary 5 and the improved lower bound in Eq. (20) to cq-channels obtaining similar results, albeit changing the lower (for the corollary) or upper (for the improvement) Minkowski dimensions of $\sqrt{\!\widetilde{\mathcal{X}}}$ to $d_{M}\!\left(\!\sqrt{\!\widehat{\mathcal{X}}}\right)$. Notice also that these are always rates for simultaneous deterministic identification because we measure the output quantum states with the fixed POVM $T^{\otimes n}$.
@@
Finally, we can also extend the results to general quantum channels $\mathcal{N}:A\rightarrow B$ with an input restriction to a subset $\mathcal{X}\subset\mathcal{S}(A)$ which can be effectively reduced to a cq-channel $W:\mathcal{X}\rightarrow\mathcal{S}(B)$, $W_{x}=\mathcal{N}(x)$. The resulting code is defined as a quantum DI $\mathcal{X}$-code (see [CDBW:DI_classical, Sect. 6]) and the corresponding lower bound on the simultaneous rate-reliability function is given, thanks to the reduction from quantum to classical-quantum channels, by Theorem 13.
@@
For the converse part, we can also follow closely the proof of Theorem 7. We have to be careful though, as in the quantum setting, we can no longer use the total variation and Euclidean distances to calculate the separation between outputs. We use instead their well-known generalization to quantum states: the trace and the Hilbert-Schmidt distances, based on the Schatten-$1$ and -$2$ norms, respectively. Given two density matrices $\rho$ and $\sigma$, these distances are related as follows [CDBW:DI_classical, Lemma 6.2]:
$\frac{1}{2}\|\rho-\sigma\|_{1}\leq\left\|\sqrt{\rho}-\sqrt{\sigma}\right\|_{2}\leq\sqrt{2}\sqrt{\|\rho-\sigma\|_{1}},$
@@
and the Fuchs-van-de-Graaf inequalities in Eq. (3) also hold [FVdG:ineq]. Notice that by conveniently defining the output quantum states as $W_{x}\in\mathcal{S}(B)$ the notation of the classical stochastic distances or fidelities between probability distributions is equivalent to the corresponding ones in the quantum setting, and we can actually follow the converse proof step by step. We fix a maximum possible fidelity between output states corresponding to code words from a good DI code, create a covering of $\sqrt{\!\widetilde{\mathcal{X}}}:=\left\{\sqrt{\rho}:\rho\in\widetilde{\mathcal{X}}\right\}\subset\mathcal{S}(B)$ such that any two states in the same ball have a fidelity higher than the initial requirement, and finally count the cardinality of such a covering. We obtain:
@@
$R(n)\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right].\qed$

### v4

@@ IV-B Zero-dimensional output probability sets
Example 12 can be generalized to DMCs with an input power constraint. Given a cost function $\phi:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ (power per symbol), an input power constraint $A$ is an upper bound on the average power that admissible code words have to obey: $\phi^{n}(x^{n})=\frac{1}{n}\sum_{t=1}^{n}\phi(x_{t})\leq A$. From [9] we know that the DI capacity of the DMC $W$ under the power constraint $A$ is
$C_{\text{DI}}(W;\phi,A)=\max_{p_{X}\,:\,\mathbb{E}\phi(X)\leq A}H(X),$ (70)
after purging the channel of duplicate input symbols: if $W_{x}=W_{x^{\prime}}$, and say $\phi(x^{\prime})\geq\phi(x)$, remove the input with the larger cost ($x^{\prime}$) from $\mathcal{X}$ and keep the other one ($x$); in case of a tie decide arbitrarily. Example 12 above is recovered when the power constraint is irrelevant (if $A\geq\max_{x\in\mathcal{X}}\phi(x)$), in which case the capacity becomes $\max_{p_{X}\in\mathcal{P}(\mathcal{X})}H(X)=\log|\mathcal{X}|=\log|W(\mathcal{X})|$. With the power constraint active, and for positive error exponents, we obtain lower and upper bounds on $R(n)$ as given in Equations (61) and (62), with $\log|W(\mathcal{X})|$ replaced by $\max_{p_{X}\,:\,\mathbb{E}\phi(X)\leq A}H(X)$.
@@
IV-C Large error exponent regime
Let us start by understanding when and why our analysis is limited to the regime of small error exponents. We have to go back to the proof of Theorem 4 where, in the second inequality of Equation (19), we use the bound $-\ln F^{2}\geq 1-F^{2}$, with $F$ the fidelity between two output distributions. Notice that this bound is only a good approximation when the letterwise fidelity is close to 1. That is, when the probability distributions are similar, which can be related to a finer packing, so smaller error exponents. Conversely, large error exponents are related to wider packings, which result on fidelity values getting closer to 0, making the bound $-\ln F^{2}\geq 1-F^{2}$ grossly inaccurate. So, when applying this bound, we limit our analysis to small error exponents.
The reader might wonder now why it is at all necessary to apply such a problematic bound. Well, in order to apply our proof idea, we need a metric so that a packing can be defined, but the negative logarithm of the fidelity [second line in Equation (17)] is not a well-defined distance, as it does not satisfy the triangle inequality. This is why we use the bound $-\ln F^{2}\geq 1-F^{2}$ together with the lower bound in Equation (9), conveniently obtaining an Euclidean metric. As we have already discussed, this step is a good approximation for the small exponent regime, which provides the framework for the superlinearity study and allows the recovery of the capacity results from [17] (which were the initial objectives of this work). However, it is increasingly inexact for bigger error exponents making the bound inadequate in the regime of large error exponents.
The lack of a lower bound able to well-approximate $-\ln F^{2}$ in the range of small fidelities (for which $-\ln F^{2}$ grows to infinity) and from which a genuine distance can be extracted, suggests that our proof will probably not be trivially generalised to the large error exponent regime. Different ideas are needed. However, we can make some comments on how the rate-reliability function behaves in the large exponent regime, as follows.
To start, we can upper-bound the rate for DI over any channel $W$ meaningfully for arbitrary error exponents starting from a transmission code for $W$. It is clear that any transmission code without randomness on the encoder and error probability $\lambda=2^{-nE_{\text{T}}}$ is also a DI code with $2^{-nE_{\text{DI}}}=\lambda_{1}=\lambda_{2}=\lambda$. This is because to identify any message $m$, we could use the transmission code to read the decoded message $\hat{m}$, and decide whether they are the same $m\stackrel{{\scriptstyle?}}{{=}}\hat{m}$, hence the DI errors and rate would then be equal to the error and the rate of the transmission code used. Thus, the rate-reliability function for transmission is a lower bound on the rate-reliability function for DI: $E_{\text{DI}}(R)\geq E_{\text{T}}(R)$. Furthermore, by using a sequential decoding of the $N=2^{nR}$ transmission messages using an $(n,N,\lambda,\lambda)$-DI code (decoding a message by trying to identify each possible output), we can bound
$2^{-nE_{\text{T}}}=\lambda\leq 2^{nR}2^{-nE_{\text{DI}}}=2^{-n(E_{\text{DI}}-R)},$ (71)
V Asymmetric error regimes
All our previous results study the performance of DI codes with errors vanishing exponentially fast: $\lambda_{1}=2^{-nE_{1}}$ and $\lambda_{2}=2^{-nE_{2}}$. The error exponents $E_{1},E_{2}>0$ are taken to be larger than some threshold $E_{1},E_{2}>E$, and this $E$ is actually the defining value of the different bounds. Therefore, the codes analysed can only produce error-symmetric asymptotic achievable triples $(R,E,E)$.
This begs the natural question: what happens to the rate when the errors are asymmetric? Can we achieve superlinear codes when only one error vanishes slowly? Here we analyse the following maximally asymmetric settings:
  • Stein regime: $\lambda_{1}<1$ and $\lambda_{2}=2^{-nE_{2}}$,
  • Sanov regime: $\lambda_{1}=2^{-nE_{1}}$ and $\lambda_{2}<1$.
These two regimes mirror two well-known classical asymmetric hypothesis testing settings. In that context, the goal is to distinguish between two probability distributions (hypotheses) $P$ and $Q$, defined on the same sample space $\mathcal{Y}$, using $n$ independent realizations of one of them. Two errors may occur: accepting $P$ (the null hypothesis) when the samples were produced by $Q$, and the converse case where $Q$ is accepted (the alternative hypothesis) while the samples come from $P$ [27].
Remark 13.
In the settings described above, we refer to the two types of errors using the parameters $\lambda_{1}$ and $\lambda_{2}$, defined explicitly in the identification context. While it may be tempting to describe these errors as “false positives” or “false negatives,” such terminology can lead to confusion, as their interpretation is not consistent across domains.
For example, in classical hypothesis testing, the null hypothesis often models a healthy patient, and the alternative an illness. A “positive” result means accepting the alternative, so a type I error is interpreted as a false positive. In identification, on the other hand, a “positive” outcome naturally refers to accepting the message under test. In this case, the first type of error (failing to identify the transmitted message) becomes a false negative, and the second (accepting an incorrect message) becomes a false positive. This reversal illustrates how informal labels, which are widely used in the literature, can obscure rather than clarify.
In the Chernoff-Stein Lemma setting, the probability $\alpha_{n}$ of accepting the alternative hypothesis when the null is true is required to be upper bounded for all $n$ by $\epsilon\in(0,1)$; then, the probability $\beta_{n}$ of accepting the null hypothesis when the alternative is true decays exponentially with $n$, and the best achievable exponent is given by the Kullback-Leibler divergence $D(P\|Q):=\sum_{y\in\mathcal{Y}}P(y)\log[P(y)/Q(y)]$ of the two distributions. Indeed,
$\lim_{n\rightarrow\infty}-\frac{1}{n}\log\beta_{n}=D(P\|Q).$ (72)
In the finite block length regime, a more refined quantity is used to characterize the tradeoff between the errors: the hypothesis testing relative entropy
$D_{h}^{\epsilon}(P\|Q):=-\log\inf_{\mathcal{S}\subset\mathcal{Y}\atop P(\mathcal{S}^{c})\leq\epsilon}Q(\mathcal{S}),$ (73)
where $\mathcal{S}$ is a decision region. This object converges to the KL divergence for increasing $n$: $\lim_{n\rightarrow\infty}\frac{1}{n}D_{h}^{\epsilon}(P^{\otimes n}\|Q^{\otimes n})=D(P\|Q)$ for any $\epsilon\in(0,1)$, connecting with the optimum exponent in the asymptotic regime.
Clearly, if we take a good DI code in the Stein regime, the identification test $\mathcal{E}_{j}$ satisfies $W_{u_{j}}(\mathcal{E}_{j}^{c})\leq\lambda_{1}$ and $W_{u_{k}}(\mathcal{E}_{j})\leq\lambda_{2}$ (for $k\neq j$). Hence we get immediately, by definition,
$D_{h}^{\lambda_{1}}(W_{u_{j}}\|W_{u_{k}})\geq-\log\lambda_{2}=nE_{2}.$ (74)
Similarly, by taking $\mathcal{E}_{k}$ as the test, we get for a good DI code in the Sanov regime that $D_{h}^{\lambda_{2}}(W_{u_{j}}\|W_{u_{k}})\geq nE_{1}$. In other words, all good DI codes in the Stein (or Sanov) regime have output distributions with a minimum pairwise hypothesis testing relative entropy linear in $n$. We will use this property to prove converses for general channels in the Stein and Sanov error settings.
We are motivated to transform the entropies above into Rényi relative entropies, which are generally easier to handle, and above all additive. Given a parameter $\alpha\in(0,\infty)\setminus\{1\}$ they are defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}}P(y)^{\alpha}Q(y)^{1-\alpha},$ (75)
and connected to the hypothesis testing relative entropies, for both $\lambda_{1}$ and $\lambda_{2}$ and $\alpha>1$ to be fixed later, through the following bound [28]:
$\!\!D_{h}^{\epsilon}(W_{u_{j}}\|W_{u_{k}})\!\leq\!D_{\alpha}(W_{u_{j}}\|W_{u_{k}})+\frac{\alpha}{1-\alpha}\log\left(\!\frac{1}{1-\epsilon}\!\right).$ (76)
As the Rényi relative entropies are additive under product distributions, we will be able to transform terms referring to code words into sums of letter-wise objects. Indeed, given two codewords $u_{j}=x_{1}\dots x_{n}\in\mathcal{X}^{n}$ and $u_{k}=x^{\prime}_{1}\dots x^{\prime}_{n}\in\mathcal{X}^{n}$, $D_{\alpha}(W_{u_{j}}\|W_{u_{k}})=\sum_{i=1}^{n}D_{\alpha}(W_{x_{i}}\|W_{x^{\prime}_{i}}).$
For our converse proof, we want to use the linear lower bounds of $D_{h}^{\lambda_{1,2}}$ described above, which in turn lower bound the sum of letter-wise Rényi entropies as follows:
$\sum_{i=1}^{n}D_{\alpha}(W_{x_{i}}\|W_{x^{\prime}_{i}})\geq nE_{1,2}-\frac{\alpha}{1-\alpha}\log\left(\frac{1}{1-\lambda_{2,1}}\right).$ (77)
Next, we create a partition of the code such that, no two different distributions in each subgroup satisfy the inequality above, meaning that each subset can only contain a single codeword. The task is therefore reduced to counting the number of subsets in the partition.
However, looking at Equation (75), we see that the sum can diverge if there are small values of $W_{x^{\prime}_{i}}(y)$ (as this term has a negative exponent in the regime $\alpha>1$). In this case, the lower bound we want to use in Eq. (77) becomes trivial. We overcome this issue by defining the following partition: let $\widetilde{\mathcal{X}}_{\ell}=W(\mathcal{X}_{\ell})$ be subsets of the output probability space $W(\mathcal{X})=:\widetilde{\mathcal{X}}=\dot{\bigcup}_{\ell=1}^{L}\widetilde{\mathcal{X}}_{\ell}$ such that for all $\ell\in[L]$, $x,x^{\prime}\in\mathcal{X}_{\ell}$, and $y\in\mathcal{Y}$:
$\frac{1}{1+\delta}W_{x^{\prime}}(y)\leq W_{x}(y)\leq(1+\delta)W_{x^{\prime}}(y),$ (78)
with $\delta>0$. With this, if for some $\ell\in[L]$ we have a problematic $W_{x^{\prime}}(y)=0$, then all the other $x\in\mathcal{X}_{\ell}$ are also $W_{x}(y)=0$, so the sum in Equation (75) will no longer diverge for distributions in the same partition (it will just be 0). In consequence, the sum in the right hand side of Eq. (77) will not diverge either. In general, for all $x_{i},x_{i}^{\prime}\in\mathcal{X}_{\ell}$
$\begin{split}D_{\alpha}(W_{x_{i}}\|W_{x^{\prime}_{i}})&=\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}}W_{x^{\prime}_{i}}(y)\left(\frac{W_{x_{i}}(y)}{W_{x^{\prime}_{i}}(y)}\right)^{\alpha}\\ &\leq\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}}W_{x^{\prime}_{i}}(y)\left(1+\delta\right)^{\alpha}\\ &=\frac{\alpha}{\alpha-1}\log(1+\delta)\stackrel{{\scriptstyle!}}{{\leq}}\frac{E_{1,2}}{2},\end{split}$ (79)
where the first inequality follows from the upper bound in Equation (78), and the last can always be true by defining the parameter $\delta>0$ suitably. In block length $n$, we will have for any two sequences $x^{n},x^{\prime n}\in\mathcal{X}_{\ell^{n}}$, where $\ell^{n}=\ell_{1}\dots\ell_{n}$ and $\mathcal{X}_{\ell^{n}}=\mathcal{X}_{\ell_{1}}\times\dots\times\mathcal{X}_{\ell_{n}}$,
$\sum_{i=1}^{n}D_{\alpha}(W_{x_{i}}\|W_{x^{\prime}_{i}})\leq n\frac{E_{1,2}}{2}.$ (80)
So, any two sequences $x^{n},x^{\prime n}\in\mathcal{X}_{\ell^{n}}$ in the same partition have maximum Rényi entropy sum given by Equation (80), and we have seen that any two codewords need to have this sum lower bounded by $nE_{1,2}-C$, with $C$ the constant last term in Equation (77). Therefore, for any big enough $n$ (specifically for any $n>n_{0}:=-2\alpha\log(1-\lambda_{1,2})/[E_{2,1}(\alpha-1)]$), each region of the partition cannot contain the outputs of two different valid codewords. I.e., necessarily, $N\leq L^{n}$. It only remains to upper-bound the size of such a partition.
Before presenting the general case, let us start with the particular family of channels which have all the output probabilities bounded away from 0 by some positive constant $\omega>0$. Notice that the partition defined above is very natural for these channels, as the problematic region (when we have probabilities close to zero) simply does not appear.
Theorem 14.
Given a channel $W$ and a constant $\omega>0$ such that for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, $W(x|y)\geq\omega$, any DI code $\mathcal{C}$ for $W$ in the Stein or Sanov regime can have at most a linear scaling rate $R(n)\leq O(1)$.
Proof.
As we have seen, we just need to upper bound the size of the partition. Given some $y\in\mathcal{Y}$ and $x\in\mathcal{X}_{\ell}$, notice that the closest probability of $y$ for an $x^{\prime}\in\mathcal{X}_{k\neq\ell}$ is at most $W_{x}(y)/(1+\delta)$ by virtue of Equation (78), and the closest probability for an $x^{\prime\prime}\in\mathcal{X}_{j\neq k,\ell}$ is $W_{x}(y)/(1+\delta)^{2}$. As a matter of fact, we will find valid $x^{(i)}$ values in $L$ different partitions until $W_{x}(y)/(1+\delta)^{L}<\omega$, as the channel does not have probabilities that small. Then, we can upper bound the number of allowed partitions $L$ with
$\omega\leq\frac{1}{(1+\delta)^{L}}\,\implies\,L\leq-\frac{\log\omega}{\log(1+\delta)}=:A(\omega,\delta).$ (81)
Where we have used that, for any pair $(x,y)$, the probability $W_{x}(y)\leq 1$ in the expression on the left, and simple algebra to find the expression on the right, where we have also defined the positive constant $A(\omega,\delta)>0$. The code size is then bounded as claimed: $|\mathcal{C}|=N\leq L^{n}\leq 2^{n\log A}$, and the rate thus $R(n)=\frac{1}{n}\log N\leq\log A(\omega,\delta)=O(1)$. ∎
This result shows that for channels with all possible output probabilities uniformly bounded away from 0, the code length scaling in the Stein or Sanov error regimes (that is, when we only force one error type to go exponentially fast to zero) can only be linear. Next, we study the general case, lifting the restriction on the minimum size of the output probabilities. We find that the superlinear scaling of order $n\log n$ (linearithmic) is excluded both for the Stein and Sanov error regimes.
Theorem 15.
Any code $\mathcal{C}$ in the Stein or Sanov regime for DI over an unrestricted channel $W$ can have at most a rate scaling as $R(n)\leq O(\log\log n)$.
Proof.
To be able to use the proof idea of Theorem 14, we have to somehow avoid the very small probability values $W_{x}(y)$. To this end, we start by defining a new channel $V$ approximating $W$ such that for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$V_{x}(y)=\left\{\frac{1}{K}W_{x}(y)\quad\text{if}\quad W_{x}(y)\geq\frac{\delta}{n|\mathcal{Y}|},\atop 0\hskip 42.67912pt\text{if}\quad W_{x}(y)<\frac{\delta}{n|\mathcal{Y}|},\right.$ (82)
with $K=\!\!\!\!\sum_{y:W_{x}(y)\geq\frac{\delta}{n|\mathcal{Y}|}}\!\!W_{x}(y)\geq 1-\frac{\delta}{n}$ a normalization constant. Here, $0<\delta<1$ is a parameter to be fixed later.
This transformed channel will have different errors $\lambda_{1}^{\prime}$ and $\lambda_{2}^{\prime}$ which we need to control in the Stein or Sanov regime. In fact, notice that the particular condition in Equation (82) has been chosen such that
$\begin{split}\frac{1}{2}\|W_{x^{n}}-V_{x^{n}}\|_{1}&\leq\frac{1}{2}\sum_{i=1}^{n}\|W_{x_{i}}-V_{x_{i}}\|_{1}\\ &=\frac{1}{2}\sum_{i=1}^{n}\sum_{y\in\mathcal{Y}}|W_{x_{i}}(y)-V_{x_{i}}(y)|\\ &\leq\frac{1}{2}\sum_{i=1}^{n}\sum_{y\in\mathcal{Y}}\frac{\delta}{n|\mathcal{Y}|}=\frac{\delta}{2},\end{split}$ (83)
where the first inequality follows from the subadditivity of the total variation distance under tensor products (triangle inequality), and the last from the bound we have chosen for our modified channel in Equation (82). Now, letting $u_{j}\neq u_{k}\in\mathcal{X}^{n}$ be two different words from a good DI code for $W$, and $\mathcal{E}_{j}\subset\mathcal{Y}^{n}$ the decision region for the word $u_{j}$, we get:
$\lambda_{1}^{\prime}$ $=\sup_{j\in[N]}V_{u_{j}}(\mathcal{E}_{j}^{c})\leq\sup_{j\in[N]}W_{u_{j}}(\mathcal{E}_{j}^{c})+\frac{\delta}{2}\leq\lambda_{1}+\frac{\delta}{2},$ (84)
$\lambda_{2}^{\prime}$ $=\!\!\!\sup_{j\neq k\in[N]}\!\!\!V_{u_{j}}(\mathcal{E}_{k})\leq\!\!\!\sup_{j\neq k\in[N]}\!\!\!W_{u_{j}}(\mathcal{E}_{k})+\frac{\delta}{2}\leq\lambda_{2}+\frac{\delta}{2}.$ (85)
When $\lambda_{1,2}<1$ and $\delta$ is small enough, the errors of the transformed channel $V$ maintain the bounded nature $\lambda^{\prime}_{1,2}<\lambda_{1,2}+\frac{\delta}{2}<1$. However, the exponential bound $\lambda_{1,2}<2^{-nE_{1,2}}$ is lost, and we have to be more careful in this regime. For this, observe the following letter-wise domination for all $x,y$: $V_{x}(y)\leq\frac{1}{1-\frac{\delta}{n}}W_{x}(y)$. Hence,
$V_{x^{n}}(y^{n})\leq\left(\frac{1}{1-\frac{\delta}{n}}\right)^{n}W_{x^{n}}(y^{n})\leq e^{2\delta}W_{x^{n}}(y^{n}),$ (86)
where we have used $\delta<1$. For any subset $\mathcal{E}\subset\mathcal{Y}^{n}$ we thus find $V_{x^{n}}(\mathcal{E})\leq e^{2\delta}W_{x^{n}}(\mathcal{E})$, so
$\lambda_{1}^{\prime}$ $=\sup_{j\in[N]}V_{u_{j}}(\mathcal{E}_{j}^{c})\leq\sup_{j\in[N]}e^{2\delta}W_{u_{j}}(\mathcal{E}_{j}^{c})\leq e^{2\delta}\lambda_{1},$ (87)
$\lambda_{2}^{\prime}$ $=\!\!\!\sup_{j\neq k\in[N]}\!\!\!V_{u_{j}}(\mathcal{E}_{k})\leq\!\!\!\sup_{j\neq k\in[N]}\!\!\!e^{2\delta}W_{u_{j}}(\mathcal{E}_{k})\leq e^{2\delta}\lambda_{2}.$ (88)
Therefore, any DI code for $V$ in the Stein error regime is a code for $W$ which, by Equations (87) and (85), is also in the Stein regime. Taking Equations (84) and (88), we can argue the same for the Sanov regime.
It only remains to upper-bound the size of a code for the transformed channel $V$. Notice that, if we restrict the input to letters with output probability distributions having the same support $\mathcal{Y}_{s}\subset\mathcal{Y}$ (with $s\in[S]$ and $S\leq 2^{|\mathcal{Y}|}$), the problem effectively becomes the one in Theorem 14, so we can reuse the partition strategy with the only difference that now the parameter $\omega$ that bounds all probabilities is a function of $n$, indeed we have imposed $\omega(n)=\frac{\delta}{n|\mathcal{Y}|}$. Directly form Equation (81) we extract:
$L\leq-\frac{\log\omega(n)}{\log(1+\delta)}=\frac{\log n-\log(\delta/|\mathcal{Y}|)}{\log(1+\delta)}=O(\log n).$ (89)
We can finally bound the size of the code by $L^{n}$ times the total number of supports in $\mathcal{Y}^{n}$, which is bounded by $|S|^{n}\leq 2^{n|\mathcal{Y}|}$:
$N\leq L^{n}2^{n|\mathcal{Y}|}=2^{n(|\mathcal{Y}|+\log L)}\leq 2^{O(n\log\log(n))},$ (90)
and therefore the rate $R(n)=\frac{1}{n}\log N\leq O(\log\log n)$. ∎
Remark 16.
The last steps in the proof above may represent a huge overcount. Indeed, it is highly improbable that different codes constructed for each support can produce a reliable general code for the overall channel without losing any word (just a sequence is received, so there is no way of knowing the support of the probability distribution that produced it). However, notice that this does not affect the size of the code at leading order, as $L$ scales with logarithmically in $n$, while the number $S$ of supports is a constant.
Furthermore, however, we do not know if the bound $R\leq O(\log\log n)$ is tight. The double logarithmic scaling could be an artifact of the proof. Indeed, we have no code constructions for the Stein or Sanov error settings that achieve such a scaling, and it is possible that $R\leq O(1)$ holds generally, as for the special case treated in Theorem 14. We leave this open problem for future investigation.
The results in this section show that having one of the two errors vanishing slowly is not enough to recover linearithmic rates, we actually need both of them to be slow. This shows, once more and in an even more dramatic manner, the fragility of the superlinear behaviour which, while being a characteristic of DI codes, appears only in a very specific, extreme error setting.
VI Extension to quantum channels
@@
Here we prove how the main results in Section III can be extended to classical-quantum (cq)-channels and to general quantum channels under the restriction that we only use product states when encoding. For the sake of brevity, we refer the reader to [17, Sect. 6] for all the necessary quantum information tools (alternatively, see [29, 30, 27]) needed for the following, as well as for a brief review on identification via quantum channels [31, 32, 33, 34, 35, 36, 37].
@@
For the direct part, we just need to define an informationally complete POVM (positive operator-valued measure) a collection of positive semidefinite operators $T=(T_{y}:y\in\mathcal{Y})$, such that $\sum_{y}T_{y}=\mathds{1}$. Then, we can form the concatenated channel $\overline{W}=\mathcal{T}\circ W$ of the cq-channel $W$ followed by the quantum-classical (qc)-channel $|\mathcal{T}(\sigma)=\sum_{y}(\operatorname{Tr}\sigma T_{y})\ket{y}\bra{y}$ |, which effectively outputs the probability of each outcome $y$ when the POVM $T$ is used as measurement. Now, since that the channel $\overline{W}$ can be identified with the classical channel $\overline{W}(y|x)=\operatorname{Tr}W_{x}T_{y}$ and $\widehat{\mathcal{X}}:=\overline{W}(\mathcal{X})\subset\mathcal{P}(\mathcal{Y})$ its image of output distributions we can reduce the problem directly to the setup of Theorem 4 to get the following:
@@
For any $0<t<1$, and $E(n)>0$, there exists a simultaneous DI code over the cq-channel $W:\mathcal{X}\rightarrow\mathcal{S}(B)$ (and its associated $\overline{W}$ as above) on block length $n$ with error exponents $E_{1}(n)\geq E(n)-\frac{1}{n}$ and $E_{2}(n)\geq E(n)-\frac{3}{n}$, and its rate lower-bounded
$R(n)$ $\geq(1-t)\log\left[\Pi_{\sqrt[4]{\frac{6E(n)}{ct^{2}}}}\left(\!\sqrt{\!\widehat{\mathcal{X}}}\right)\right]$ (91)
$\phantom{\geq.}-H(t,1-t)-O\left(\frac{\log n}{n}\right).\qed$
@@
With the same arguments, we can also extend Corollary 6 and the improved lower bound in Equation (39) to cq-channels obtaining similar results, albeit changing the lower (for the corollary) or upper (for the improvement) Minkowski dimension of $\sqrt{\!\widetilde{\mathcal{X}}}$ to that of $\sqrt{\!\widehat{\mathcal{X}}}$. We remark that these are always rates for simultaneous deterministic identification because we measure the output quantum states with the fixed POVM $T^{\otimes n}$. See [38] for a non-simultaneous code construction that can achieve higher rates in terms of $d_{M}\!\left(\!\!\sqrt{\!\widetilde{\mathcal{X}}}\right)$.
@@
We can also extend the results to general quantum channels $\mathcal{N}:A\rightarrow B$ with an input restriction to a subset $\mathcal{X}\subset\mathcal{S}(A)$ which can be effectively reduced to a cq-channel $W:\mathcal{X}\rightarrow\mathcal{S}(B)$, $W_{x}=\mathcal{N}(x)$. The resulting code is defined as a quantum DI $\mathcal{X}$-code (see [17, Sect. 6]) and the corresponding lower bound on the simultaneous rate-reliability function is given, thanks to the reduction from quantum to classical-quantum channels, by Theorem 17.
@@
For the converse, we can follow closely the proof of Theorem 7. We have to be careful, though, as in the quantum setting we can no longer use the total variation and Euclidean distances to calculate the separation between outputs. We use instead their well-known generalization to quantum states: the trace and the Hilbert-Schmidt distances, based on the Schatten-111 and -222 norms, respectively. Given two density matrices $\rho$ and $\sigma$, these distances are related as follows [17, Lemma 6.2]:
$\frac{1}{2}\|\rho-\sigma\|_{1}\leq\left\|\sqrt{\rho}-\sqrt{\sigma}\right\|_{2}\leq\sqrt{2}\sqrt{\|\rho-\sigma\|_{1}},$ (92)
@@
and the Fuchs-van-de-Graaf inequalities in Equation (8) also hold [39]. Notice that by denoting the output quantum states as $W_{x}\in\mathcal{S}(B)$, the notation of the classical stochastic distances or fidelities between probability distributions is equivalent to the corresponding ones in the quantum setting, and we can actually follow the converse proof step by step. We fix a maximum possible fidelity between output states corresponding to code words from a good DI code, create a covering of $\sqrt{\!\widetilde{\mathcal{X}}}:=\left\{\sqrt{\rho}:\rho\in\widetilde{\mathcal{X}}\right\}\subset\mathcal{S}(B)$ such that any two states in the same ball have a fidelity higher than the initial requirement, and finally count the cardinality of such a covering. We obtain:
@@
$R(n)\leq\log\left[\Gamma_{\frac{1}{2}{\sqrt{1-e^{-E(n)/2}}}}\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\right].\qed$ (93)